2.) 20J of work are required to move 2.0C of charge between two points in an electric field. a.)Find the electric potential difference between the two points.




3.)How much work would be required to move 5.0C of charge between the same two points of the problem above?




Please answer with a detailed explanation & Solution.



Thanks

1
Expert's answer
2022-04-07T13:09:03-0400

(а)

"V=\frac{W}{q}=\frac{20\:\rm J}{2.0\:\rm C}=10\:\rm V"

(b)

"W=qV=10*5.0=50\:\rm J"